OpenStudy (anonymous):

The exterior angles of an octagon are 42°, 55°, 39°, 20°, 62°, 45°, and 47°. What is the measure of the eighth exterior angle? Show equations and all work that leads to your answer. Help Please(:

OpenStudy (anonymous):

@jigglypuff314 Help?

jigglypuff314 (jigglypuff314):

for any polygon, the sum of all it's exterior angles is 360 degrees http://www.mathsisfun.com/geometry/exterior-angles-polygons.html so 42 + 55 + 39 + 20 + 62 + 45 + 47 + x = 360 add like terms then solve for x :)

OpenStudy (anonymous):

So 42 + 55 + 39 + 20 + 62 + 45 + 47=310 so x must equal fifty to =360?

jigglypuff314 (jigglypuff314):

umm... solve for x when given x + 310 = 360

jigglypuff314 (jigglypuff314):

oh yeah nvm x = 50 is right :)

OpenStudy (anonymous):

x=50 then right

OpenStudy (anonymous):

Thankkks

jigglypuff314 (jigglypuff314):

you're welcome :)